# suppReferences

### 17 Mar 2024 v1

@@ Introduction
Finally, we found that the quality of Muse-generated corrections to misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Methods; Fig. 3b; Supplementary Fig. S7-S11). The quality of Muse-generated corrections is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated corrections. The quality of Muse-generated corrections is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness corrections by laypeople.
@@ Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[21, 22], our work demonstrates the potential of LLMs to improve the online information ecosystem by correcting misinformed, inaccurate, and misleading information[2, 3]. Correcting misinformation, particularly on social media, is complex. It requires comprehending multimodal content, understanding context involving emerging topics or events, accurately and comprehensively identifying stated inaccuracies that are subtly different from the facts, and writing easy-to-understand, precise, and informative explanations supported by trustworthy references. We propose Muse and demonstrate the high quality of its automatically generated corrections to misinformation on social media. Muse augments existing powerful LLMs (e.g., GPT-4[26]) with the capabilities of addressing visuals, accessing up-to-date knowledge, and finding accurate references nonparametrically (i.e., without the need to train or fine-tune a machine learning model) such that Muse can be easily and cheaply updated, especially compared to parametric models[35]. Results further validate that addressing visual information is crucial in correcting multimodal misinformation, where GPT-4 struggles but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining inaccuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4.
@@
We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently correcting misinformation. Muse is end-to-end and thereby simple to use. Meanwhile, our results reveal that Muse-generated corrections have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness corrections in correcting misinformation across modalities and domains (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated corrections show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating corrections with the least toxic text and most credible references[31, 41, 42] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[25] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality corrections by Muse can generally be obtained two minutes after misinformation appears on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify corrections themselves (Fig 2,3f).
@@
Muse’s corrections cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the complexity of this task and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[18] to identify whether a new article’s headline and lede contains misinformation without explaining the inaccuracies.
@@
This work also faces the following limitations. First, although Muse is capable of addressing and correcting multimodal misinformation with text and images, it cannot accept video inputs. Second, we focus on correcting misinformation in a single language, English, one of the most spoken languages in the world, whereas misinformation has become a global issue and exists in various high- and low-resource languages [45]. Third, we evaluated Muse using real-world misinformation on a single social media platform, X, as its Community Notes system provides publicly accessible large-scale data and a transparent state-of-the-art aggregation algorithm, enabling us to compare Muse’s corrections with the high- and average-helpfulness corrections made by social media users. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4[26], which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
@@ Conclusion
We proposed Muse, a nonparametric LLM for scaling up high-quality and prompt corrections of misinformation on social media, that can integrate textual and visual information, access timely updated knowledge, and generate easy-to-understand corrections with accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness corrections on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of corrections overall. Muse also excels in correcting text and multimodal (textual and visual) misinformation and misinformation across politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. This study demonstrates the potential of LLMs in correcting misinformation online effectively, scalably, promptly, and transparently.
@@ Acknowledgements
We would like to thank Hacks/Hackers for advertising the study and helping with the recruitment. We also thank members of the UW Social Futures Lab and Behavioral Data Science Group for their suggestions and feedback. This work was supported in part by the National Science Foundation’s Convergence Accelerator program under Award No. 49100421C0037.
@@
X.Z., A.X.Z., and T.A. designed the study. X.Z. led, and A.S. assisted in developing the model. A.S. built the web interface for evaluation. X.Z. prepared and analyzed the data. All authors contributed to writing the manuscript. A.X.Z. and T.A. supervised the study.
@@
Figure S1: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each need to be corrected, whereas generating one query may overlook some misinformed, inaccurate, or misleading claims. Bold text: the misinformed, inaccurate, or misleading claims overlooked when generating one query but captured when generating more than one query.
Figure S2: Examples that show how retrieved web pages with relatively low relevance to misinformation can promote LLM (in this case, GPT-4) hallucinations when generating corrections.
@@
Figure S4: An example of a high-helpfulness correction from Community Notes displayed on the corresponding tweet on X and visible to the public.
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S5: Expert evaluation results for text misinformation (#=155).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S6: Expert evaluation results for multimodal (textual and visual) misinformation (#=77).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S7: Expert evaluation results for political and international-affair misinformation (#=80).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S8: Expert evaluation results for economic and business misinformation (#=38).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S9: Expert evaluation results for crime and law misinformation (#=38).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S10: Expert evaluation results for social-issue and human-right misinformation (#=30).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S11: Expert evaluation results for health and medicine misinformation (#=24).
@@
Figure S14: Distribution of helpfulness scores of laypeople’s corrections in X Community Notes. All: All laypeople’s corrections in Community Notes. Average helpfulness: Laypeople’s corrections in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s corrections in Community Notes identified with high helpfulness and used in our study. For $x\pm y$, $x$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
Figure S15: Distribution of creation times of laypeople’s corrections in X Community Notes used in our study. Median of the creation time of laypeople’s average-helpfulness corrections: 16 hours after the misinformation was posted. Median of the creation time of laypeople’s high-helpfulness corrections: 13 hours after the misinformation was posted.
@@
Figure S24: Impact of starting times of correcting misinformation on Muse’s performance. The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness correction was created (median: 16 hours after the corresponding misinformation was posted; Supplementary Fig. S15). The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness correction was created (median: 13 hours after the corresponding misinformation was posted; Supplementary Fig. S15). The simulated starting time for Muse: The post time of the corresponding misinformation (i.e., 0 hours after the corresponding misinformation was posted).
Figure S25: Distribution of post times of misinformation in X Community Notes included in our analyses.
Figure S26: Impact of post times of misinformation on Muse’s performance.
(a) Overall quality of corrections.
(b) Quality of corrections in identifying and explaining inaccuracies.
(c) Quality of corrections in generated text.
(d) Quality of corrections in references.

### 06 Apr 2024 v2

@@ Introduction
Finally, we found that the quality of Muse-generated responses to potential misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Methods; Fig. 3b; Supplementary Fig. S9-S13). The quality of Muse-generated responses is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated responses. The quality of Muse-generated responses is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness responses by laypeople.
@@ Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[42, 43], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[1, 2]. Real-world misinformation can combine (in)accurate, factually accurate but misleading, and unverifiable claims. Identifying a piece of content as misinformation without further interventions, done by many previous AI-driven approaches[44], has limited impact on reducing misinformation’s detrimental effects. Showing the identification results without explanations might even increase false beliefs[45]. Correcting misinformation has been shown to reduce its spread and false beliefs, whose effectiveness is affected by quality, timeliness, and scalability[4, 22]. However, research on scaling up and accelerating misinformation correction—which may include automation—is still at the early stage. Existing AI models struggle to identify and correct misinformation, especially on social media, where misinformation is not restricted to narrow domains and spreads rapidly. Correcting misinformation requires comprehending content that can be multimodal and the context beyond it that often involves emerging events. It requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content is (in)accurate with trustworthy references. We propose Muse and demonstrate the high quality of its automatically generated responses to social media posts that potentially are misinformation. Muse augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4.
@@
We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently responding to suspicious content. Muse is end-to-end and thereby simple to use. It is nonparametric, i.e., does not need training or fine-tuning an AI model, and thereby easily and cheaply updated, especially compared to parametric models[32]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities and domains (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[28, 38, 39] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[22] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality responses by Muse can generally be obtained two minutes after suspicious content (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f).
@@
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation.
@@
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
@@ Conclusion
We proposed Muse, a nonparametric LLM to scale up and accelerate misinformation corrections on social media. Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness responses on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of responses overall. Muse also excels when the responded content is textual or multimodal (textual and visual) and when it is related to politics and international affairs, economy and business, crime and law, social issues and human rights, or health and medicine. This study demonstrates the potential of LLMs in responding to online misinformation effectively, scalably, promptly, and transparently.
@@ Acknowledgements
We would like to thank Hacks/Hackers for advertising the study and helping with the recruitment. We also thank members of the UW Social Futures Lab and Behavioral Data Science Group for their suggestions and feedback. This work was supported in part by the National Science Foundation’s Convergence Accelerator program under Award No. 49100421C0037. T.A. and A.S. were supported in part by the Office of Naval Research (#N00014-21-1-2154), NSF IIS-1901386, and NSF CAREER IIS-2142794.
@@
X.Z., A.X.Z., and T.A. designed the study. X.Z. led, and A.S. assisted in developing the model. A.S. built the web interface for evaluation. X.Z. prepared and analyzed the data. X.Z. drafted and all authors revised the manuscript. A.X.Z. and T.A. supervised the study.
@@
Figure S1: Distribution of potential misinformation in X Community Notes (as of February 2023) that received its first response (gray) or first high-quality response (orange) within a certain amount of time.
Figure S2: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each needs to be verified, whereas generating one query may overlook some of them and hence lead to not comprehensive identifications of (in)accuracies. Bold text: the verification-needed claims that are overlooked when generating one query but captured when generating more than one query.
Figure S3: Examples that show how retrieved web pages with relatively low relevance to potential misinformation can promote LLM (in this case, GPT-4) hallucinations when generating responses.
@@
Figure S5: Examples of Muse-generated responses to accurate, partially accurate, and factually correct but misleading content on social media.
Figure S6: An example of a high-helpfulness response from Community Notes displayed on the corresponding tweet and visible to the public.
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S7: Expert evaluation results for textual posts (#=155).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S8: Expert evaluation results for multimodal (textual and visual) posts (#=77).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S9: Expert evaluation results for political and international-affair posts (#=80).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S10: Expert evaluation results for economic and business posts (#=38).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S11: Expert evaluation results for crime and law posts (#=38).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S12: Expert evaluation results for social-issue and human-right posts (#=30).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S13: Expert evaluation results for health and medicine posts (#=24).
@@
Figure S16: Distribution of helpfulness scores of laypeople’s responses in X Community Notes. All: All laypeople’s responses in Community Notes. Average helpfulness: Laypeople’s responses in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study. For $x\pm y$, $x$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
Figure S17: Distribution of creation times of laypeople’s responses in X Community Notes used in our study. Median of the creation time of laypeople’s average-helpfulness responses: 16 hours after the tweet was posted. Median of the creation time of laypeople’s high-helpfulness responses: 13 hours after the tweet was posted.
@@
Figure S26: Impact of starting times of responding to tweets on Muse’s performance. The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness responses was created (median: 16 hours after the corresponding tweet was posted; Supplementary Fig. S17). The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness responses was created (median: 13 hours after the corresponding tweet was posted; Supplementary Fig. S17). The simulated starting time for Muse: The post time of the corresponding tweet (i.e., 0 hours after the corresponding tweet was posted).
Figure S27: Distribution of post times of tweets in X Community Notes included in our analyses.
Figure S28: Impact of post times of tweets on Muse’s performance.
(a) Overall quality of responses.
(b) Quality of responses in identifying and explaining (in)accuracies.
(c) Quality of responses in generated text.
(d) Quality of responses in references.